# Refereed Learning

### v1

@@ 3.1 Certifiable sample and certifiable sum
To prove Item 1, we will argue that the distributions $\widehat{{\mathcal{D}}}^{\prime}$ and $\Phi_{\delta}{\left[{{\mathcal{D}}^{\prime}}\right]}$ (Definition 3.3) are equivalent. Then, we will argue that sampling $i\sim{\mathcal{D}}^{\prime}$ and outputting $x\leftarrow S_{i}$ yields the same distribution as sampling $x\sim{\mathcal{D}}_{\lambda}$. Finally, we will complete the proof by leveraging the guarantee of Birgé’s decomposition for monotone distributions (3.4), and by arguing the ${\mathcal{D}}_{\lambda}$ and $\mathcal{D}$} are close in total variation distance. First, by Lemma 3.2, the verifier correctly obtains $T_{\lambda}=\sum_{x\in\{0,1\}^{d}}\lfloor{{\mathcal{D}}(x)}\rfloor_{\lambda}$, and thus can provide itself with query access to $Q_{{\mathcal{D}}_{\lambda}}$ by first querying $Q_{{\mathcal{D}}}(x)$ to obtain ${\mathcal{D}}(x)$ and then computing $Q_{{\mathcal{D}}_{\lambda}}(x)=\frac{\lfloor{{\mathcal{D}}(x)}\rfloor_{\lambda}}{T_{\lambda}}$.
@@
If either ${\mathcal{P}}_{0}$ or ${\mathcal{P}}_{1}$ is honest, then the distribution $\widehat{{\mathcal{D}}}^{\prime}$ constructed by $\mathcal{V}$ in Protocol 1 is equivalent to the flattened histogram $\Phi_{\delta}{\left[{{\mathcal{D}}^{\prime}}\right]}$ defined in Definition 3.3.
@@
Assume without loss of generality that ${\mathcal{P}}_{0}$ is honest. Then by Lemma 3.2, the verifier obtains the correct value of $N=|S|$. Similarly, by 3.5, for all $k\in[\ell]$ the verifier obtains the correct endpoints $L_{k},R_{k}\in S$ of the set $S[I_{k}]$. Given the endpoints of $S[I_{k}]$ and query access to the probability mass function $Q_{{\mathcal{D}}_{\lambda}}$ of distribution ${\mathcal{D}}_{\lambda}$, the verifier can provide itself with query access to $t(x)={\mathcal{D}}_{\lambda}(x)\cdot\mathds{1}{\left[{x\in S[I_{k}]}\right]}$. Thus, by Lemma 3.2, the verifier obtains the correct values of $p_{k}={\mathcal{D}}_{\lambda}(S[I_{k}])$ for each $k\in[\ell]$. It follows that for all $k\in[\ell]$ and $i\in[k]$ we have
$\widehat{{\mathcal{D}}}^{\prime}(i)=\frac{{\mathcal{D}}_{\lambda}(S[I_{k}])}{|I_{k}|}=\frac{{\mathcal{D}}^{\prime}(I_{k})}{|I_{k}|}=\Phi_{\delta}{\left[{{\mathcal{D}}^{\prime}}\right]}(i).\qed$
By definition of $\mathcal{D}$$^{\prime}$, sampling $j\sim{\mathcal{D}}^{\prime}$ and outputting $x\leftarrow S_{j}$ is equivalent to sampling $x\sim{\mathcal{D}}_{\lambda}$. Moreover, since $\mathcal{D}$$^{\prime}$ is a monotone distribution over $[N]$, 3.4 immediately implies that $d_{\mathrm{TV}}({\mathcal{D}}^{\prime},\Phi_{\delta}{\left[{{\mathcal{D}}^{\prime}}\right]})\leq\delta$. Let $\widehat{{\mathcal{D}}}_{\lambda}$ be the distribution given by sampling $j\sim\widehat{{\mathcal{D}}}^{\prime}$ and outputting $x\leftarrow S_{j}$. By 3.6, we have $d_{\mathrm{TV}}({\mathcal{D}}^{\prime},\widehat{{\mathcal{D}}}^{\prime})\leq\delta$ and hence $d_{\mathrm{TV}}({\mathcal{D}}_{\lambda},\widehat{{\mathcal{D}}}_{\lambda})\leq\delta$ as well. To complete the proof of Item 1, it remains to argue that $d_{\mathrm{TV}}({\mathcal{D}}_{\lambda},{\mathcal{D}})\leq\delta$, and hence $d_{\mathrm{TV}}(\widehat{{\mathcal{D}}}_{\lambda},{\mathcal{D}})\leq 2\delta$.
Claim 3.7.
Fix $d,\lambda\in\mathbb{N}$ with $\lambda>d$ and a distribution $\mathcal{D}$} over $\{0,1\}^{d}$. Let ${\mathcal{D}}_{\lambda}$ be the distribution defined by ${\mathcal{D}}_{\lambda}(x)=\frac{\lfloor{{\mathcal{D}}(x)}\rfloor_{\lambda}}{\sum_{x\in\{0,1\}^{d}}\lfloor{{\mathcal{D}}(x)}\rfloor_{\lambda}}$, where $\lfloor{y}\rfloor_{\lambda}$ denotes $2^{-\lambda}\cdot\lfloor{2^{\lambda}\cdot y}\rfloor$ for all $y\in\mathbb{R}$. Then ${\mathcal{D}}_{\lambda}$ is $\lambda$-precise and $d_{\mathrm{TV}}({\mathcal{D}}_{\lambda},{\mathcal{D}})\leq 2^{d+1-\lambda}$.
Proof.
Observe that for each $y\in[0,1]$ we have ${\left|{y-\lfloor{y}\rfloor_{\lambda}}\right|}\leq 2^{-\lambda}$, and therefore ${\left|{\lfloor{{\mathcal{D}}(x)}\rfloor_{\lambda}-{\mathcal{D}}(x)}\right|}\leq 2^{-\lambda}$ for each $x\in\{0,1\}^{d}$. It follows that $\sum_{x\in\{0,1\}^{d}}{\left|{\lfloor{{\mathcal{D}}(x)}\rfloor_{\lambda}-{\mathcal{D}}(x)}\right|}\leq 2^{d-\lambda}$ and hence, since $\sum_{x\in\{0,1\}^{d}}{\mathcal{D}}(x)=1$, that $\sum_{x\in\{0,1\}^{d}}\lfloor{{\mathcal{D}}(x)}\rfloor_{\lambda}\in{\left[{1\pm 2^{d-\lambda}}\right]}$. Let $T_{\lambda}=\sum_{x\in\{0,1\}^{d}}\lfloor{{\mathcal{D}}(x)}\rfloor_{\lambda}$. Then $T_{\lambda}\in{\left[{1\pm 2^{d-\lambda}}\right]}$. Let $a\in[\pm 2^{d-\lambda}]$ be such that $T_{\lambda}=1+a$. Then
@@
$\leq\frac{2^{d-\lambda}}{1-2^{d-\lambda}}\leq 2^{d+1-\lambda}.$
To see that ${\mathcal{D}}_{\lambda}$ is $\lambda$-precise, observe that ${\mathcal{D}}_{\lambda}(x)=\frac{\lfloor{2^{\lambda}{\mathcal{D}}(x)}\rfloor}{\sum\lfloor{2^{\lambda}{\mathcal{D}}(x)}\rfloor}$ has both numerator and denominator that are non-negative integers at most $2^{\lambda}$. ∎
By 3.7 and our choice of $\lambda$ in Protocol 1, we have $d_{\mathrm{TV}}(\widehat{{\mathcal{D}}}_{\lambda},{\mathcal{D}})\leq 2\delta$, and hence setting $\delta\leftarrow\delta/2$ completes the proof of Item 1. To see why Item 2 holds, recall that by Lemmas 3.2 and 3.5, each call to certifiable sum with $\lambda=d+\log\frac{4+\delta}{\delta}$ uses $\lambda\mathop{\mathrm{poly}}d=\log\frac{1}{\delta}\cdot\mathop{\mathrm{poly}}d$ bits of communication and verifier runtime, and each call to certifiable index, uses a verifier that runs in time $\mathop{\mathrm{poly}}(d)$ and uses $\mathop{\mathrm{poly}}(d)$ bits of communication. Since $N\leq 2^{d}$, and certifiable sum is called once for each of the $\ell=\Theta{\left({\frac{\log N}{\delta}}\right)}$ partitions, computing the distribution $\widehat{{\mathcal{D}}}^{\prime}$ takes the verifier $\frac{1}{\delta}\log\frac{1}{\delta}\cdot\mathop{\mathrm{poly}}d$ time, and uses $\frac{1}{\delta}\log\frac{1}{\delta}\cdot\mathop{\mathrm{poly}}d$ bits of communication. Similarly, since the verifier draws $m$ samples and calls certifiable index for each sample, the total runtime of the verifier and communication complexity of the protocol is ${\left({\frac{1}{\delta}\log\frac{1}{\delta}+m}\right)}\cdot\mathop{\mathrm{poly}}d$. ∎
In the remainder of the section we complete the proof of our second tool, the certifiable sum protocol (Lemma 3.2), and leverage it to complete the proof of the certifiable index protocol (3.5). At a high level, the protocol works as follows: For all $b\in\{0,1\}$ let $C_{b}={\left\{{x\in\{0,1\}^{d}\mid x_{1}=b}\right\}}$. First, ${\mathcal{P}}_{0}$ claims that the sum over $\{0,1\}^{d}$ is $\widehat{T}$, and that the sum over $C_{0}$ and $C_{1}$ is $\widehat{T}_{0}$ and $\widehat{T}_{1}$ respectively. Since $C_{0}$ and $C_{1}$ are disjoint and $C_{0}\cup C_{1}=\{0,1\}^{d}$, we must have $\widehat{T}_{0}+\widehat{T}_{1}=\widehat{T}$. The key observation is that if $\widehat{T}\neq T$ (the correct value of the sum), then either $\widehat{T}_{0}\neq T_{0}$ or $\widehat{T}_{1}\neq T_{1}$ (the correct values of the sum on $C_{0}$ and $C_{1}$). The verifier can then ask ${\mathcal{P}}_{1}$ for the bit $b$ such that $\widehat{T}_{b}\neq T_{b}$, and repeat the above steps on $C_{b}$. After $d$ rounds there is only a single point remaining in the subcube, and hence the verifier can check if the claim made by ${\mathcal{P}}_{0}$ is correct by making a single query to $t$. The protocol to certify a claim made by ${\mathcal{P}}_{1}$ is identical but has the roles of ${\mathcal{P}}_{0}$ and ${\mathcal{P}}_{1}$ switched.
@@
Our protocol consists of two phases. In the first phase the verifier uses ${\mathcal{P}}_{0}$ to verify the claim made by ${\mathcal{P}}_{1}$, and in the second phase the verifier uses ${\mathcal{P}}_{1}$ to verify the claim made by ${\mathcal{P}}_{0}$. The final protocol executes both phases and returns the first verified claim (this will be correct if at least one prover is honest). We first define the following notation: for all $t:\{0,1\}^{d}\to\mathbb{Q}_{\lambda}$ and points $z\in\{0,1\}^{j}$ for some $j<d$, let
@@
To see why the communication complexity holds, observe that the numerator in $\sum_{x\in\{0,1\}^{d}}t(x)$ can be at most $2^{d+2\lambda}$, and that the denominator can be at most $2^{\lambda\cdot d}$. Thus, $\widehat{T}_{z}$ will require at most $O(d\lambda)$ bits to send. Since the protocol uses $2d$ rounds, and in each round $O(d\lambda)$ bits are required to transmit $(\widehat{T}_{z},\widehat{T}_{z0},\widehat{T}_{z1})$, it immediately follows that the overall communication complexity is $\lambda\mathop{\mathrm{poly}}d$. The runtime and query complexity of the verifier follows by inspection of Protocol 2. ∎
Proof of 3.5.
List of Protocols 3 certifiable index
Without loss of generality assume ${\mathcal{P}}_{0}$ is honest. Then $\widehat{x}=S_{i}$ and hence |A|=i1|A|=i-1. By Lemma 3.2, protocol ${\left[{{\mathcal{P}}_{0},{\mathcal{P}}^{*}_{1},\mathcal{V}}\right]}_{0}$ outputs accept. Next, consider ${\left[{{\mathcal{P}}_{0},{\mathcal{P}}^{*}_{1},\mathcal{V}}\right]}_{b}$. If ${\mathcal{P}}^{*}_{1}$ sends $\widehat{x}\neq S_{i}$, then, if $\widehat{x}\not\in S$, the protocol rejects in Step 2. On the other hand, if $\widehat{x}\in S$ then $|A|=|{\left\{{x\in S\mid x\prec\widehat{x}}\right\}}\neq i-1$. By Lemma 3.2, the sum $\sum_{x\in\{0,1\}^{d}}t(x)=|A|\neq i-1$, and hence the protocol outputs reject. To complete the proof, define the protocol $\left[{{\mathcal{P}}_{0},{\mathcal{P}}_{1},\mathcal{V}}\right]$} as follows: for each $b\in\{0,1\}$ run protocols ${\left[{{\mathcal{P}}_{0},{\mathcal{P}}_{1},\mathcal{V}}\right]}_{b}$, and return the $\widehat{x}$ from the first round of an accepting execution. By the above argument, the protocol $\left[{{\mathcal{P}}_{0},{\mathcal{P}}_{1},\mathcal{V}}\right]$} always outputs $S_{i}$. The communication complexity and runtime guarantees follow by inspection of Protocol 3 and from the guarantees of Lemma 3.2.∎
@@ 3.2 Refereed query delegation
Fix a deterministic oracle $\mathcal{O}$. Suppose there exists a protocol $\left[{{\mathcal{P}}_{0},{\mathcal{P}}_{1},\mathcal{V}}\right]$} that for all inputs $\kappa\in\mathbb{R}$ and $\lambda\in\mathbb{N}$ has communication complexity $C(\kappa,\lambda)$, verifier runtime $T_{\mathcal{V}}(\kappa,\lambda)$, verifier query complexity $q_{\mathcal{V}}(\kappa,\lambda)$, prover runtime $T_{\mathcal{P}}(\kappa,\lambda)$, and prover query complexity $q_{\mathcal{P}}(\kappa,\lambda)$. Additionally, assume all queries to $\mathcal{O}$ and their answers can be specified using at most $\lambda$ bits.
Then there exists a protocol $\left[{\widetilde{{\mathcal{P}}}_{0},\widetilde{{\mathcal{P}}}_{1},\widetilde{\mathcal{V}}}\right$]} such that for all $b\in\{0,1\}$ and $\widetilde{{\mathcal{P}}}^{*}_{1-b}$ there exists a ${\mathcal{P}}^{*}_{1-b}$ such that
Moreover, $\left[{\widetilde{{\mathcal{P}}}_{0},\widetilde{{\mathcal{P}}}_{1},\widetilde{\mathcal{V}}}\right$]} has communication complexity $C(\kappa,\lambda)+2\lambda\cdot q_{\mathcal{V}}(\kappa,\lambda)$, verifier runtime $T_{\mathcal{V}}(\kappa,\lambda)$, prover runtime $T_{\mathcal{P}}(\kappa,\lambda)+q_{\mathcal{V}}(\kappa,\lambda)$, prover query complexity $q_{\mathcal{P}}(\kappa,\lambda)+q_{\mathcal{V}}(\kappa,\lambda)$, and verifier query complexity at most $1$.
Proof.
@@
    ${\left[{\widetilde{{\mathcal{P}}}^{\mathcal{O}}_{0},\widetilde{{\mathcal{P}}}^{\mathcal{O}}_{1},\widetilde{\mathcal{V}}^{\mathcal{O}}}\right]}(\kappa,\lambda)$ 1. Simulate ${\mathcal{P}}_{0}^{\mathcal{O}},{\mathcal{P}}_{1}^{\mathcal{O}},\mathcal{V}^{\mathcal{O}}](\kappa,\lambda)$, except answer $\mathcal{V}$’s queries to $\mathcal{O}$ using the following procedure: (a) $\widetilde{\mathcal{V}}$ asks both $\widetilde{{\mathcal{P}}}_{0}$ and $\widetilde{{\mathcal{P}}}_{1}$ to answer the query using their access to $\mathcal{O}$. (b) If both provers agree, continue the protocol using this query answer. (c) If the provers disagree, $\widetilde{\mathcal{V}}$ makes a single query to $\mathcal{O}$ to determine the lying prover (1b)(1-b), and then uses the answers from $\widetilde{{\mathcal{P}}}_{b}$ for all subsequent queries to $\mathcal{O}$.
List of Protocols 4 refereed query delegation
The communication complexity, runtime, and query complexity follow immediately from the fact that Protocol 4 simply runs ${\mathcal{P}}_{0},{\mathcal{P}}_{1},\mathcal{V}]$ once, and has each prover make at most $q_{\mathcal{V}}(\kappa,\lambda)$ additional queries to $\mathcal{O}$.
@@
Let ${\mathcal{P}}^{*}_{1-b}$ be the prover that is identical to $\widetilde{{\mathcal{P}}}^{*}_{1-b}$ (except it simulates query requests from $\mathcal{V}$ and does not actually send query answers to $\mathcal{V}$). [^8] Morally, think of ${\mathcal{P}}^{*}_{1-b}$ and $\widetilde{{\mathcal{P}}}^{*}_{1-b}$ as identical. However, a subtlety arises that the behavior of $\widetilde{{\mathcal{P}}}^{*}_{1-b}$ may depend on the queries it answers for $\mathcal{V}$—e.g., $\widetilde{{\mathcal{P}}}^{*}_{1-b}$ may execute strategy “a” when asked an oracle query by $\mathcal{V}$, and strategy “b” when $\mathcal{V}$ does not ask any oracle queries. To ensure that ${\mathcal{P}}^{*}_{1-b}$ executes the correct malicious prover strategy, it simulates $\widetilde{{\mathcal{P}}}^{*}_{1-b}$ with the queries from $\mathcal{V}$. We see that, if $\mathcal{V}$ is simulated using queries to $\mathcal{O}$ that are answered correctly, then Protocol 4 has exactly the same distribution of outputs as ${\left[{{\mathcal{P}}^{\mathcal{O}}_{b},{\mathcal{P}}^{*}_{1-b},\mathcal{V}^{\mathcal{O}}}\right]}(\kappa,\lambda)$.

### v2

@@ 3.1 Certifiable sample and certifiable sum
To prove the lemma, it suffices to show (1) that each loop of the protocol obtains an i.i.d. sample $\widehat{x}\sim{\mathcal{D}}_{\lambda}$ and (2) that $d_{\mathrm{TV}}({\mathcal{D}},{\mathcal{D}}_{\lambda})\leq\delta$.
We first show item (1). Note that, for fixed $p$, there is a unique value $\widehat{x}_{b}$, and this a unique value $p^{\mathit{small}}_{b}$, satisfying the conditions; let $x^{*}$ and $p^{*}$ denote these values. Assume without loss of generality that ${\mathcal{P}}_{0}$ is honest. The value $\widehat{x}_{0}$ is accepted: by the correctness guarantees of certifiable sum (Lemma 3.2), since ${\mathcal{P}}_{0}$ is honest then $\mathcal{V}$ correctly obtains $\widehat{x}=x^{*}$ and $p^{\mathit{small}}_{0}=p^{*}$, so $\widehat{x}_{0}$ will satisfy the condition checked by $\mathcal{V}$ on Line 3(b)iii. Next note that the value $\widehat{x}_{1}$ is rejected if $\widehat{x}_{1}\neq x^{*}$, since in this case it fails to satisfy the condition $p^{\mathit{small}}_{1}<p\leq p^{\mathit{small}}_{1}+\lfloor{{\mathcal{D}}(\widehat{x}_{1})}\rfloor_{\lambda}.$ Thus, by the correctness guarantees of certifiable sum (Lemma 3.2), and since ${\mathcal{P}}_{0}$ is honest, the verifier obtains a value $\widehat{x}=x^{*}$ satisfying the condition checked on Line 3(b)iii.
We now prove that if $\widehat{x}$ satisfies the condition $p_{\mathit{small}}<p\leq p_{\mathit{small}}+\lfloor{{\mathcal{D}}(\widehat{x})}\rfloor_{\lambda}$, then $\widehat{x}\sim{\mathcal{D}}_{\lambda}$. This follows immediately from
$\operatorname*{{Pr}}[\widehat{x}=x]$ $=\operatorname*{{Pr}}_{p}{\left[{\sum_{x^{\prime}\prec x}\lfloor{{\mathcal{D}}(x^{\prime})}\rfloor_{\lambda}<p\leq\sum_{x^{\prime}\preceq x}\lfloor{{\mathcal{D}}(x^{\prime})}\rfloor_{\lambda}}\right]}$
$=\frac{\sum_{x^{\prime}\preceq x}\lfloor{{\mathcal{D}}(x^{\prime})}\rfloor_{\lambda}-\sum_{x^{\prime}\prec x}\lfloor{{\mathcal{D}}(x^{\prime})}\rfloor_{\lambda}}{\sum_{x\in\{0,1\}^{d}}\lfloor{{\mathcal{D}}(x)}\rfloor_{\lambda}}$
$=\frac{\lfloor{{\mathcal{D}}(x)}\rfloor_{\lambda}}{\sum_{x\in\{0,1\}^{d}}\lfloor{{\mathcal{D}}(x)}\rfloor_{\lambda}}={\mathcal{D}}_{\lambda}(x).$
Because each of the $m$ iterations of the protocol uses an i.i.d. value $p\sim{\mathcal{D}}_{\lambda}$, we obtain an i.i.d. value $\widehat{x}\sim{\mathcal{D}}_{\lambda}$ from each iteration of the protocol.
To complete the proof of Item 1, we argue that $d_{\mathrm{TV}}({\mathcal{D}}_{\lambda},{\mathcal{D}})\leq\delta$.
@@
Fix $d,\lambda\in\mathbb{N}$ with $\lambda>d$ and a distribution $\mathcal{D}$} over $\{0,1\}^{d}$. Let ${\mathcal{D}}_{\lambda}$ be the distribution defined by ${\mathcal{D}}_{\lambda}(x)=\frac{\lfloor{{\mathcal{D}}(x)}\rfloor_{\lambda}}{\sum_{x\in\{0,1\}^{d}}\lfloor{{\mathcal{D}}(x)}\rfloor_{\lambda}}$, where $\lfloor{y}\rfloor_{\lambda}$ denotes $2^{-\lambda}\cdot\lfloor{2^{\lambda}\cdot y}\rfloor$ for all $y\in\mathbb{R}$. Then $d_{\mathrm{TV}}({\mathcal{D}}_{\lambda},{\mathcal{D}})\leq 2^{d+1-\lambda}$.
@@
Observe that for each $y\in[0,1]$ we have ${\left|{y-\lfloor{y}\rfloor_{\lambda}}\right|}\leq 2^{-\lambda}$, and therefore ${\left|{\lfloor{{\mathcal{D}}(x)}\rfloor_{\lambda}-{\mathcal{D}}(x)}\right|}\leq 2^{-\lambda}$ for each $x\in\{0,1\}^{d}$. It follows that $\sum_{x\in\{0,1\}^{d}}{\left|{\lfloor{{\mathcal{D}}(x)}\rfloor_{\lambda}-{\mathcal{D}}(x)}\right|}\leq 2^{d-\lambda}$ and hence, since $\sum_{x\in\{0,1\}^{d}}{\mathcal{D}}(x)=1$, that $\sum_{x\in\{0,1\}^{d}}\lfloor{{\mathcal{D}}(x)}\rfloor_{\lambda}\in{\left[{1\pm 2^{d-\lambda}}\right]}$. Thus, for $T_{\lambda}=\sum_{x\in\{0,1\}^{d}}\lfloor{{\mathcal{D}}(x)}\rfloor_{\lambda}$ we have $T_{\lambda}\in{\left[{1\pm 2^{d-\lambda}}\right]}$. Let $a\in[\pm 2^{d-\lambda}]$ be such that $T_{\lambda}=1+a$. Then
@@
$\leq\frac{2^{d-\lambda}}{1-2^{d-\lambda}}\leq 2^{d+1-\lambda},$
completing the proof. ∎
To see why the communication complexity holds, observe the protocol consists of four calls to certifiable sum and the communication of a constant number of additional terms with $\lambda=d+1+\log\frac{1}{\delta}$ bits of precision. Thus, up to constant factors the runtime and communication complexity are inherited from certifiable sum. ∎
In the remainder of the section we complete the proof of our second tool, the certifiable sum protocol (Lemma 3.2). At a high level, the protocol works as follows: For all $b\in\{0,1\}$ let $C_{b}={\left\{{x\in\{0,1\}^{d}\mid x_{1}=b}\right\}}$. First, ${\mathcal{P}}_{0}$ claims that the sum over $\{0,1\}^{d}$ is $\widehat{T}$, and that the sum over $C_{0}$ and $C_{1}$ is $\widehat{T}_{0}$ and $\widehat{T}_{1}$ respectively. Since $C_{0}$ and $C_{1}$ are disjoint and $C_{0}\cup C_{1}=\{0,1\}^{d}$, we must have $\widehat{T}_{0}+\widehat{T}_{1}=\widehat{T}$. The key observation is that if $\widehat{T}\neq T$ (the correct value of the sum), then either $\widehat{T}_{0}\neq T_{0}$ or $\widehat{T}_{1}\neq T_{1}$ (the correct values of the sum on $C_{0}$ and $C_{1}$). The verifier can then ask ${\mathcal{P}}_{1}$ for the bit $b$ such that $\widehat{T}_{b}\neq T_{b}$, and repeat the above steps on $C_{b}$. After $d$ rounds there is only a single point remaining in the subcube, and hence the verifier can check if the claim made by ${\mathcal{P}}_{0}$ is correct by making a single query to $t$. The protocol to certify a claim made by ${\mathcal{P}}_{1}$ is identical but has the roles of ${\mathcal{P}}_{0}$ and ${\mathcal{P}}_{1}$ switched.
@@
Our protocol consists of two phases. In the first phase the verifier uses ${\mathcal{P}}_{0}$ to verify the claim made by ${\mathcal{P}}_{1}$, and in the second phase the verifier uses ${\mathcal{P}}_{1}$ to verify the claim made by ${\mathcal{P}}_{0}$. The final protocol executes both phases and returns the first verified claim (this will be correct if at least one prover is honest). We first define the following notation: for all $t:\{0,1\}^{d}\to\mathsf{Prec}_{\lambda}$ and points $z\in\{0,1\}^{j}$ for some $j<d$, let
@@
To see why the communication complexity holds, observe that sum $\sum_{x\in\{0,1\}^{d}}t(x)$ is bounded above by $2^{d+2\lambda}$ and is an integer multiple of $2^{-\lambda}$. Thus, $\widehat{T}_{z}$ will require at most $O(d\lambda)$ bits to send. Since the protocol uses $2d$ rounds, and in each round $O(d\lambda)$ bits are required to transmit $(\widehat{T}_{z},\widehat{T}_{z0},\widehat{T}_{z1})$, it immediately follows that the overall communication complexity is $\lambda\mathop{\mathrm{poly}}d$. The runtime and query complexity of the verifier follows by inspection of Protocol 2. ∎
@@ 3.2 Refereed query delegation
Fix a deterministic oracle $\mathcal{O}$. Let $\left[{{\mathcal{P}}_{0},{\mathcal{P}}_{1},\mathcal{V}}\right]$} be a protocol that, for all inputs $\kappa\in\mathbb{R}$ and $\lambda\in\mathbb{N}$, has communication complexity $C(\kappa,\lambda)$, verifier runtime $T_{\mathcal{V}}(\kappa,\lambda)$, verifier query complexity $q_{\mathcal{V}}(\kappa,\lambda)$, prover runtime $T_{\mathcal{P}}(\kappa,\lambda)$, and prover query complexity $q_{\mathcal{P}}(\kappa,\lambda)$. Additionally, assume all queries to $\mathcal{O}$ and their answers can be specified using at most $\lambda$ bits.
Then for all $b\in\{0,1\}$ and $\widetilde{{\mathcal{P}}}^{*}_{1-b}$ there exists a ${\mathcal{P}}^{*}_{1-b}$ such that
${\left[{\widetilde{{\mathcal{P}}}^{\mathcal{O}}_{b},\widetilde{{\mathcal{P}}}^{*}_{1-b},\widetilde{\mathcal{V}}^{\mathcal{O}}}\right]}(\kappa,\lambda)={\left[{{\mathcal{P}}^{\mathcal{O}}_{b},{\mathcal{P}}^{*}_{1-b},\mathcal{V}^{\mathcal{O}}}\right]}(\kappa,\lambda),$
where $\left[{\widetilde{{\mathcal{P}}}_{0},\widetilde{{\mathcal{P}}}_{1},\widetilde{\mathcal{V}}}\right$]} are described in Protocol 3. Moreover, $\left[{\widetilde{{\mathcal{P}}}_{0},\widetilde{{\mathcal{P}}}_{1},\widetilde{\mathcal{V}}}\right$]} has communication complexity $C(\kappa,\lambda)+2\lambda\cdot q_{\mathcal{V}}(\kappa,\lambda)$, verifier runtime $T_{\mathcal{V}}(\kappa,\lambda)$, prover runtime $T_{\mathcal{P}}(\kappa,\lambda)+q_{\mathcal{V}}(\kappa,\lambda)$, prover query complexity $q_{\mathcal{P}}(\kappa,\lambda)+q_{\mathcal{V}}(\kappa,\lambda)$, and verifier query complexity at most $1$.
Proof of Lemma 3.4.
@@
    ${\left[{\widetilde{{\mathcal{P}}}^{\mathcal{O}}_{0},\widetilde{{\mathcal{P}}}^{\mathcal{O}}_{1},\widetilde{\mathcal{V}}^{\mathcal{O}}}\right]}(\kappa,\lambda)$ 1. Simulate ${\mathcal{P}}_{0}^{\mathcal{O}},{\mathcal{P}}_{1}^{\mathcal{O}},\mathcal{V}^{\mathcal{O}}](\kappa,\lambda)$, except answer $\mathcal{V}$’s queries to $\mathcal{O}$ using the following procedure: (a) $\widetilde{\mathcal{V}}$ asks both $\widetilde{{\mathcal{P}}}_{0}$ and $\widetilde{{\mathcal{P}}}_{1}$ to answer the query using their access to $\mathcal{O}$. (b) If both provers agree, continue the protocol using this query answer. (c) If the provers disagree, $\widetilde{\mathcal{V}}$ makes a single query to $\mathcal{O}$ to determine the lying prover (1b)(1-b), and then uses the answers from $\widetilde{{\mathcal{P}}}_{b}$ for all subsequent queries to $\mathcal{O}$. [^7] For the competing prover setting: At the cost of one additional query, the verifier can catch a malicious $\widetilde{{\mathcal{P}}}_{b}$ in this phase by asking $\widetilde{{\mathcal{P}}}_{1-b}$ which of $\widetilde{{\mathcal{P}}}_{b}$’s queries are incorrect at the end of the protocol.
List of Protocols 3 refereed query delegation
The communication complexity, runtime, and query complexity follow immediately from the fact that Protocol 3 simply runs ${\mathcal{P}}_{0},{\mathcal{P}}_{1},\mathcal{V}]$ once, and has each prover make at most $q_{\mathcal{V}}(\kappa,\lambda)$ additional queries to $\mathcal{O}$.
@@
Let ${\mathcal{P}}^{*}_{1-b}$ be the prover that is identical to $\widetilde{{\mathcal{P}}}^{*}_{1-b}$ (except it simulates query requests from $\mathcal{V}$ and does not actually send query answers to $\mathcal{V}$). [^8] Morally, think of ${\mathcal{P}}^{*}_{1-b}$ and $\widetilde{{\mathcal{P}}}^{*}_{1-b}$ as identical. However, a subtlety arises that the behavior of $\widetilde{{\mathcal{P}}}^{*}_{1-b}$ may depend on the queries it answers for $\mathcal{V}$—e.g., $\widetilde{{\mathcal{P}}}^{*}_{1-b}$ may execute strategy “a” when asked an oracle query by $\mathcal{V}$, and strategy “b” when $\mathcal{V}$ does not ask any oracle queries. To ensure that ${\mathcal{P}}^{*}_{1-b}$ executes the correct malicious prover strategy, it simulates $\widetilde{{\mathcal{P}}}^{*}_{1-b}$ with the queries from $\mathcal{V}$. We see that, if $\mathcal{V}$ is simulated using queries to $\mathcal{O}$ that are answered correctly, then Protocol 3 has exactly the same distribution of outputs as ${\left[{{\mathcal{P}}^{\mathcal{O}}_{b},{\mathcal{P}}^{*}_{1-b},\mathcal{V}^{\mathcal{O}}}\right]}(\kappa,\lambda)$.